# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 3.1. Preliminaries
Here, update (3.10a) enforces the consensus constraint $A\mathbf{x}=0$ via an ascent step; (3.10b) forms a momentum/correction term using past dual/primal states; (3.10c) performs a proximal gradient-type step on each block $x_{i}$ whereas (3.10d) is a projected ascent ensuring $\theta_{i}^{k+1}\in\mathcal{K}_{i}^{\ast}$. At this point, we have an abstract algorithm as described in (3.10) with momentum acceleration in $\mathbf{x}$-updates to exploit the curvature of $f(\mathbf{x})$ and $G(\mathbf{x})$, and it is in a suitable form to design a backtracking condition for it. Next, we discuss how this abstract update rule can be implemented in distributed manner using only local communications for $n$-dimensional data vectors.
@@ 3.2. Distributed implementation of the abstract method
Therefore, (3.10) can be implemented in a distributed manner, i.e., each node $i\in\mathcal{N}$, starting from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$, initializes $s_{i}^{0}=s_{i}^{-1}=0$, $x_{i}^{-1}=x_{i}^{0}$ and $\theta_{i}^{-1}=\theta_{i}^{0}$, sets $r_{i}^{-1}=\mathbf{J}g_{i}(x_{i}^{0})^{\!\top}\theta_{i}^{0}$, and computes the following updates: [^4] Our initialization implies that $p_{i}^{0}=r_{i}^{0}=\mathbf{J}g_{i}(x_{i}^{0})^{\!\top}\theta_{i}^{0}$ for all $i\in\mathcal{N}$.
@@
${\rm prox}_{\tau\phi_{i}}(x)\triangleq\arg\min_{w\in\mathbb{R}^{n}}\Big\{\tau\phi_{i}(w)+\frac{1}{2}\|w-x\|^{2}\Big\}.$ (3.12)
Thus, each node $i\in\mathcal{N}$ updates its local variables $(x_{i},\theta_{i},s_{i})$ using only local gradients/Jacobians and communicating only with the neighboring nodes in $\mathcal{N}_{i}$.
The accelerated primal-dual algorithm with backtracking (ADPB) introduced in [hamedani2021primal] is a version of ADP that supports backtracking; however it was designed for centralized minimax problems. Our goal in this paper is to design a distributed stepsize selection mechanism for the abstract method given in (3.11) so that we can extend ADPB [hamedani2021primal] from centralized to the decentralized setting. The distributed stepsize selection procedure we propose is based on local backtracking and it does not require a-priori knowledge of global Lipschitz constants.
@@ 3.3. From the abstract method towards D-APDB in algorithm 1
Here we propose a step-size choice mechanism for the abstract distributed update scheme in (3.11), which would lead to D-APDB displayed in Algorithm 1. Our goal in each iteration $k\geq 0$ is to employ an Armijo-type local backtracking condition for determining local step-sizes $\tau_{i}^{k},\sigma_{i}^{k}>0$ for each $i\in\mathcal{N}$, and to use a global-max across the network to coordinate the momentum parameter $\eta^{k}>0$.
@@
$\tilde{p}_{i}$ $\leftarrow r_{i}+\frac{\tau_{i}^{-}}{\tilde{\tau}_{i}}(r_{i}-r_{i}^{-}),$ (3.13)
$\tilde{x}_{i}$ $\leftarrow\operatorname{prox}_{\tilde{\tau}_{i}\phi_{i}}\!\Big(x_{i}-\tilde{\tau}_{i}\big(\nabla f_{i}(x_{i})+\tilde{p}_{i}\big)\Big),$ (3.14)
$\tilde{\theta}_{i}$ $\leftarrow\mathcal{P}_{K_{i}^{*}\cap\mathcal{B}_{i}}\!\big(\theta_{i}+\tilde{\sigma}_{i}\,g(\tilde{x}_{i})\big),$ (3.15)
where $\rho\in(0,1)$ is a fixed contraction factor, $\delta\in(0,1)$ is a backtracking parameter, and $\tilde{p}_{i}$ denotes a local momentum term mimicking the update in (3.11c). We adopted an Armijo-type backtracking condition based on a local merit function $E_{i}$ as defined in eq. 2.4, where for given trial primal stepsize $\tilde{\tau}_{i}$, the parameters within $E_{i}$ depends only on $\tilde{\tau}_{i}$ and $\tau_{i}^{-}$, i.e.,
$\tilde{\alpha}_{i}^{k+1}=\frac{c_{\alpha}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\beta}_{i}^{k+1}=\frac{c_{\beta}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\varsigma}_{i}^{k+1}=\frac{c_{\varsigma}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\sigma}_{i}^{k}=\zeta_{i}\tilde{\tau}_{i}^{k},\quad\eta_{i}^{k}=\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}},$ (3.16)
@@
$\eta^{k}\;=\;\max_{i\in\mathcal{N}}\eta_{i}^{k}.$ (3.18)
This step requires a global max consensus, i.e., the cooperation of the nodes to compute a maximum over the node variables $\eta_{i}^{k}$. It can be typically implemented with protocols such as LoRaWAN (low power, long range) which supports long-distance low-power wireless communication between the nodes [chen2025parameter, kim2016low]. Then, D-APDB derives $\tau_{i}^{k}={\tau_{i}^{k-1}}/{\eta^{k}},\ \alpha_{i}^{k+1}=c_{\alpha}/\tau_{i}^{k},\ \beta_{i}^{k+1}=c_{\beta}/\tau_{i}^{k},\ \varsigma_{i}^{k+1}=c_{\varsigma}/\tau_{i}^{k},\ \sigma_{i}^{k+1}=\zeta_{i}\tau_{i}^{k},$ and chooses $\gamma_{k}$ via
$\gamma_{k}\;=\;\frac{c_{\gamma}}{\bar{\tau}}(\frac{2}{c_{\alpha}}+\frac{\eta^{k}}{c_{\varsigma}})^{-1}\!,$ (3.19)
with $\bar{\tau}\triangleq\max_{i\in\mathcal{N}}\{\bar{\tau}_{i}\}$ and for some $c_{\gamma}>0$ such that $c_{\gamma}\leq 1/(2|\mathcal{E}|)$, to update the $s_{i}^{k}$ variable. With this information at hand, the subsequent updates consist of updating the node-based $p_{i}^{k}$, $r_{i}^{k}$ and $x_{i}^{k}$ variables as outlined in Algorithm 1. Basically, these updates correspond to a fully decentralized implementation of (3.11) –more precisely, Algorithm 1 describes how primal-dual stepsize sequence $\{\tau_{i}^{k},\sigma_{i}^{k}\}_{k\geq 0}$ for $i\in\mathcal{N}$ and the momentum paraneter sequence $\{\eta^{k}\}_{k\geq 0}$ defining the updates rule in (3.11) can be chosen through employing node-specific backtracking and running one time max-consensus across the network per iteration $k\geq 0$.
@@ 4. Preliminary Technical Results
Now, we are ready to give our main technical results. We will derive some key inequalities below for D-APDB iterates $\{\mathbf{x}_{k},\lambda_{k},\theta_{k}\}_{k\geq 0}$ generated by Algorithm 1. Let $\mathbf{y}=[\theta^{\top},\lambda^{\top}]^{\top}$ denote the concatenation of the dual variables and $\Phi:\mathop{\bf dom}\phi\times\mathcal{K}^{*}\times\mathbb{R}^{n|\mathcal{E}|}\to\mathbb{R}$ denote the coupling function:
$\Phi(\mathbf{x},\theta,\lambda)=\langle G(\mathbf{x}),\theta\rangle+\langle A\mathbf{x},\lambda\rangle.$ (4.1)
Thus, we have $\mathcal{L}(\mathbf{x},\theta,\lambda)=\varphi(\mathbf{x})+\Phi(\mathbf{x},\theta,\lambda)-h(\theta)$, where $\mathcal{L}$ and $h(\cdot)$ as defined in (3.1), and $\varphi(\cdot)$ as defined in (2.1). For $k\geq 0$, define
@@
$q_{i}^{k,\mathbf{x}}$ $\triangleq\nabla_{x_{i}}\Phi(\mathbf{x}^{k},\theta^{k-1},\lambda^{k-1})-\nabla_{x_{i}}\Phi(\mathbf{x}^{k-1},\theta^{k-1},\lambda^{k-1}).$
@@
for any $\alpha_{i}^{k},\beta_{i}^{k}>0$. We will use eq. 4.4 in the later proof a few times.
@@
Suppose Assumptions 2, 3 and 1 hold. Fix $k\geq 0$ and let $\mathbf{x}^{k},\mathbf{x}^{k-1}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{\bf dom}\phi}$, $\theta^{k},\theta^{k-1}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{\bf dom}h}$ and $\lambda^{k},\lambda^{k-1}\in\mathbb{R}^{n|\mathcal{E}|}$ be arbitrary. Let $\{\tau_{i}^{k}\}_{i\in\mathcal{N}}$, $\{\sigma_{i}^{k}\}_{i\in\mathcal{N}}$, $\eta^{k}$ and $\gamma^{k}$ be arbitrary positive scalars, and suppose $(\mathbf{x}^{k+1},\theta^{k+1},\lambda^{k+1})$ is generated from $(\mathbf{x}^{k},\theta^{k},\lambda^{k})$ according to (3.10).
For any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathop{\bf dom}h$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, let $\mathbf{z}\triangleq(\mathbf{x},\mathbf{y})$ with $\mathbf{y}\triangleq(\theta,\lambda)$, and similarly $\mathbf{y}^{k+1}\triangleq(\theta^{k+1},\lambda^{k+1})$. Then, for any collection of positive parameters $\{\alpha_{i}^{k},\alpha_{i}^{k+1},\beta_{i}^{k},\beta_{i}^{k+1},\varsigma_{i}^{k}$, $\varsigma_{i}^{k+1}\}_{i\in\mathcal{N}}$, the following inequality holds:
@@
Fix arbitrary $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathop{\bf dom}h$ and $\lambda$. Using [hamedani2021primal, Lemma 7.1] for the $\mathbf{x}$-, $\theta$- and $\lambda$- subproblems in (3.10), we get
@@
By combining LABEL:eq:y_concave and 4.7b, we can have
@@
On the other hand, combining eqs. 4.7a and 4.8, and adding both $f(\mathbf{x}^{k+1})$ and $\Phi(\mathbf{x},\theta^{k+1},\lambda^{k+1})$ to both sides, for $k\geq 0$, we get
@@
where the inequality follows from eq. 4.7c. Thus, the above inequalities with eq. 4.12 imply
@@
Thus, given some arbitrary $\alpha_{i}^{k},\beta_{i}^{k}>0$ for $i\in\mathcal{N}$ and $k\geq 0$, eq. 4.4 implies that
@@
$\quad+\frac{1}{2\gamma^{k}}\|\lambda-\lambda^{k+1}\|^{2}+\langle q^{k+1},\mathbf{x}-\mathbf{x}^{k+1}\rangle+\langle A\mathbf{x}^{k+1}-A\mathbf{x}^{k},\lambda^{k+1}-\lambda\rangle+\sum_{i\in\mathcal{N}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\varsigma_{i}^{k+1}}{2}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}\Big]$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\Lambda^{k}}+\sum_{i\in\mathcal{N}}\Big(\frac{1}{2\alpha_{i}^{k+1}}\|q_{i}^{k+1,\mathbf{y}}\|^{2}+\frac{1}{2\beta_{i}^{k+1}}\|q_{i}^{k+1,\mathbf{x}}\|^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\varsigma_{i}^{k+1}}{2}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\eta^{k}}{2\varsigma_{i}^{k}}d_{i}}\|\lambda^{k+1}-\lambda^{k}\|^{2}\Big),$
@@
Given some arbitrary initial points $\mathbf{x}^{0}\in\mathop{\bf dom}\phi$, $\theta^{0}\in\mathcal{K}^{*}$, $\lambda^{0}\in\mathbb{R}^{n|\mathcal{E}|}$, step sizes $\gamma^{k}>0$, $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$, and the momentum parameter $\eta^{k}>0$ for all $k\geq 0$, we set $\mathbf{x}^{-1}=\mathbf{x}^{0}$, $\theta^{-1}=\theta^{0}$, $\lambda^{-1}=\lambda^{0}$, and consider the iterations as in (3.10). Suppose there exists $\{t_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$ such that the algorithm parameters satisfy the following inequalities:
@@
Next, we argue that under the parameter rule in (4.14), for $k=1$ to K2K-2, we have $t_{k+1}Q^{k+1}(\mathbf{z})-t_{k}R^{k+1}(\mathbf{z})\leq 0$; indeed, it is easy to verify after writing this inequality equivalently as follows using the definitions of $Q^{k}(\mathbf{z})$ and $R^{k+1}(\mathbf{z})$ given in (4.6a) and (4.6b), respectively:
@@
Given arbitrary $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}\in\mathop{\bf dom}\phi$ and $\theta^{\prime},\theta^{\prime\prime}\in\mathcal{K}^{*}\cap\mathcal{B}$. Recall that $\Phi(\mathbf{x},\theta,\lambda)=\langle G(\mathbf{x}),\theta\rangle+\langle A\mathbf{x},\lambda\rangle$ for any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$. Thus, for any $i\in\mathcal{N}$ and $\lambda$, we have

### v2

@@ 3.1. Preliminaries
Here, update (3.8a) enforces the consensus constraint $A\mathbf{x}=0$ via an ascent step; (3.8b) forms a momentum/correction term using past dual/primal states; (3.8c) performs a proximal gradient-type step on each block $x_{i}$ whereas (3.8d) is a projected ascent ensuring $\theta_{i}^{k+1}\in\mathcal{K}_{i}^{\ast}$. At this point, we have an abstract algorithm as described in (3.8) with momentum acceleration in $\mathbf{x}$-updates to exploit the curvature of $G(\mathbf{x})$, and it is in a suitable form to design a backtracking condition for it. Next, we discuss how this abstract update rule can be implemented in a distributed manner using only local communications for $n$-dimensional data vectors.
@@ 3.2. Distributed implementation of the abstract method
Therefore, (3.8) can be implemented in a distributed manner, i.e., each node $i\in\mathcal{N}$, starting from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$, initializes $s_{i}^{0}=s_{i}^{-1}=0$, $x_{i}^{-1}=x_{i}^{0}$ and $\theta_{i}^{-1}=\theta_{i}^{0}$, sets $r_{i}^{-1}=\mathbf{J}g_{i}(x_{i}^{0})^{\!\top}\theta_{i}^{0}$, and computes the following updates4:4Our initialization implies that $p_{i}^{0}=r_{i}^{0}=\mathbf{J}g_{i}(x_{i}^{0})^{\!\top}\theta_{i}^{0}$ for all $i\in\mathcal{N}$.
@@
${\rm prox}_{\tau\phi_{i}}(x)\triangleq\arg\min_{w\in\mathbb{R}^{n}}\Big\{\tau\phi_{i}(w)+\frac{1}{2}\|w-x\|^{2}\Big\}.$
Thus each node $i\in\mathcal{N}$ updates its local variables $(x_{i},\theta_{i},s_{i})$ using only local gradients/Jacobians and communicating only with the neighboring nodes in $\mathcal{N}_{i}$.
The accelerated primal-dual algorithm with backtracking (ADPB) introduced in [20] is a version of ADP that supports backtracking; however it was designed for centralized minimax problems. Our goal in this paper is to design a distributed stepsize selection mechanism for the abstract method given in (3.9) so that we can extend ADPB [20] from centralized to the decentralized setting. The distributed stepsize selection procedure we propose is based on local backtracking and it does not require a priori knowledge of global Lipschitz constants.
@@ 3.3. From the abstract method towards D-APDB in Algorithm 1
Here we propose a step-size choice mechanism for the abstract distributed update scheme in (3.9), which would lead to D-APDB displayed in Algorithm 1. Our goal in each iteration $k\geq 0$ is to employ an Armijo-type local backtracking condition for determining local step-sizes $\tau_{i}^{k},\sigma_{i}^{k}>0$ for each $i\in\mathcal{N}$, and to use a global-max consensus across the network to coordinate the momentum parameter $\eta^{k}>0$.
@@
$\tilde{p}_{i}$ $\leftarrow r_{i}+\frac{\tau_{i}^{-}}{\tilde{\tau}_{i}}(r_{i}-r_{i}^{-}),$
$\tilde{x}_{i}$ $\leftarrow\operatorname{prox}_{\tilde{\tau}_{i}\phi_{i}}\!\Big(x_{i}-\tilde{\tau}_{i}\big(\nabla f_{i}(x_{i})+\tilde{p}_{i}\big)\Big),$
$\tilde{\theta}_{i}$ $\leftarrow\mathcal{P}_{K_{i}^{*}\cap\mathcal{B}_{i}}\!\big(\theta_{i}+\tilde{\sigma}_{i}\,g_{i}(\tilde{x}_{i})\big),$
where $\rho\in(0,1)$ is a fixed contraction factor, $\delta\in(0,1)$ is a backtracking parameter, and $\tilde{p}_{i}$ denotes a local momentum term mimicking the update in (3.9c). We adopted an Armijo-type backtracking condition based on a local merit function $E_{i}$ as defined in (2.2), where for given trial primal stepsize $\tilde{\tau}_{i}$, the parameters within $E_{i}$ depend only on $\tilde{\tau}_{i}$ and $\tau_{i}^{-}$, i.e.,
$\tilde{\alpha}_{i}^{k+1}=\frac{c_{\alpha}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\beta}_{i}^{k+1}=\frac{c_{\beta}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\varsigma}_{i}^{k+1}=\frac{c_{\varsigma}}{\tilde{\tau}_{i}^{k}},\quad\tilde{\sigma}_{i}^{k}=\zeta_{i}\tilde{\tau}_{i}^{k},\quad\eta_{i}^{k}=\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}},$
@@
$\eta^{k}\;=\;\max_{i\in\mathcal{N}}\eta_{i}^{k}.$
This step requires a global max consensus, i.e., the cooperation of the nodes to compute a maximum over the node variables $\eta_{i}^{k}$. It can be typically implemented with protocols such as LoRaWAN (low power, long range) which supports long-distance low-power wireless communication between the nodes [14, 25]. Then, D-APDB derives $\tau_{i}^{k}={\tau_{i}^{k-1}}/{\eta^{k}},\ \alpha_{i}^{k+1}=c_{\alpha}/\tau_{i}^{k},\ \beta_{i}^{k+1}=c_{\beta}/\tau_{i}^{k},\ \varsigma_{i}^{k+1}=c_{\varsigma}/\tau_{i}^{k},\ \sigma_{i}^{k+1}=\zeta_{i}\tau_{i}^{k},$ and chooses $\gamma_{k}$ via
$\gamma_{k}\;=\;\frac{c_{\gamma}}{\bar{\tau}}(\frac{2}{c_{\alpha}}+\frac{\eta^{k}}{c_{\varsigma}})^{-1}\!,$
with $\bar{\tau}\triangleq\max_{i\in\mathcal{N}}\{\bar{\tau}_{i}\}$ and for some $c_{\gamma}>0$ such that $c_{\gamma}\leq 1/(2|\mathcal{E}|)$, to update the $s_{i}^{k}$ variable. With this information at hand, the subsequent updates consist of updating the node-based $p_{i}^{k}$, $r_{i}^{k}$ and $x_{i}^{k}$ variables as outlined in Algorithm 1. Basically, these updates correspond to a fully decentralized implementation of (3.9) –more precisely, Algorithm 1 describes how primal-dual stepsize sequence $\{\tau_{i}^{k},\sigma_{i}^{k}\}_{k\geq 0}$ for $i\in\mathcal{N}$ and the momentum parameter sequence $\{\eta^{k}\}_{k\geq 0}$ defining the update rule in (3.9) can be chosen through employing node-specific backtracking and running the max-consensus protocol across the network once per iteration $k\geq 0$.
@@ 4. Preliminary Technical Results
Now, we are ready to give our main technical results. We will derive some key inequalities below for D-APDB iterates $\{\mathbf{x}_{k},\lambda_{k},\theta_{k}\}_{k\geq 0}$ generated by Algorithm 1. Let $\mathbf{y}=[\theta^{\top},\lambda^{\top}]^{\top}$ denote the concatenation of the dual variables and $\Phi:\mathop{\bf dom}\phi\times\mathcal{K}^{*}\times\mathbb{R}^{n|\mathcal{E}|}\to\mathbb{R}$ denote the coupling function: $\Phi(\mathbf{x},\theta,\lambda)=\langle G(\mathbf{x}),\theta\rangle+\langle A\mathbf{x},\lambda\rangle.$ Thus, we have $\mathcal{L}(\mathbf{x},\theta,\lambda)=\varphi(\mathbf{x})+\Phi(\mathbf{x},\theta,\lambda)-h(\theta)$, where $\mathcal{L}$ and $h(\cdot)$ as defined in (3.1), and $\varphi(\cdot)$ as defined in (2.1). For $k\geq 0$, define
@@
$q_{i}^{k,\mathbf{x}}$ $\triangleq\nabla_{x_{i}}\Phi(\mathbf{x}^{k},\theta^{k-1},\lambda^{k-1})-\nabla_{x_{i}}\Phi(\mathbf{x}^{k-1},\theta^{k-1},\lambda^{k-1})$
@@
for any $\alpha_{i}^{k},\beta_{i}^{k}>0$. We will use (4.3) later in the proof a few times.
Remark 4.1.
For $i\in\mathcal{N}$ such that $L_{g_{i}}=0$, for all $k\geq 0$, one trivially has $q_{i}^{k,\mathbf{x}}=0;$ hence, setting $\beta_{i}^{k}=0,$ we adopt $0^{2}/0=0$ and set $\|q_{i}^{k,\mathbf{x}}\|^{2}/\beta_{i}^{k}=0.$
@@
Suppose that Assumptions 2, 3 and 1 hold. Fix $k\geq 0$ and let $\mathbf{x}^{k},\mathbf{x}^{k-1}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{\bf dom}\phi}$, $\theta^{k},\theta^{k-1}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{\bf dom}h}$ and $\lambda^{k},\lambda^{k-1}\in\mathbb{R}^{n|\mathcal{E}|}$ be arbitrary. Let $\{\tau_{i}^{k}\}_{i\in\mathcal{N}}$, $\{\sigma_{i}^{k}\}_{i\in\mathcal{N}}$, $\eta^{k}$ and $\gamma^{k}$ be arbitrary positive scalars, and suppose $(\mathbf{x}^{k+1},\theta^{k+1},\lambda^{k+1})$ is generated from $(\mathbf{x}^{k},\theta^{k},\lambda^{k})$ according to (3.8). For any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathop{\bf dom}h$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, let $\mathbf{z}\triangleq(\mathbf{x},\mathbf{y})$ with $\mathbf{y}\triangleq(\theta,\lambda)$, and similarly $\mathbf{y}^{k+1}\triangleq(\theta^{k+1},\lambda^{k+1})$. Then, for any collection of positive parameters $\{\alpha_{i}^{k},\alpha_{i}^{k+1},\beta_{i}^{k},\beta_{i}^{k+1},\varsigma_{i}^{k}$, $\varsigma_{i}^{k+1}\}_{i\in\mathcal{N}}$, the following inequality holds:
@@
Fix arbitrary $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathop{\bf dom}h$ and $\lambda$. Using [20, Lemma 7.1] for the $\mathbf{x}$-, $\theta$- and $\lambda$- subproblems in (3.8), we get
@@
By combining (LABEL:eq:y_concave) and (4.6b), we have
@@
On the other hand, by combining Equations (4.6a) and (4.7), and adding both $f(\mathbf{x}^{k+1})$ and $\Phi(\mathbf{x},\theta^{k+1},\lambda^{k+1})$ to both sides, for $k\geq 0$, we get
@@
where the inequality follows from (4.6c). Thus, the above inequalities with (4.11) implies
@@
Thus, given some arbitrary $\alpha_{i}^{k},\beta_{i}^{k}>0$ for $i\in\mathcal{N}$ and $k\geq 0$, (4.3) implies that
@@
$\quad+\frac{1}{2\gamma^{k}}\|\lambda-\lambda^{k+1}\|^{2}+\langle q^{k+1},\mathbf{x}-\mathbf{x}^{k+1}\rangle+\langle A\mathbf{x}^{k+1}-A\mathbf{x}^{k},\lambda^{k+1}-\lambda\rangle$
$\quad+\sum_{i\in\mathcal{N}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\varsigma_{i}^{k+1}}{2}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}\Big]$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\Lambda^{k}}+\sum_{i\in\mathcal{N}}\Big(\frac{1}{2\alpha_{i}^{k+1}}\|q_{i}^{k+1,\mathbf{y}}\|^{2}+\frac{1}{2\beta_{i}^{k+1}}\|q_{i}^{k+1,\mathbf{x}}\|^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\varsigma_{i}^{k+1}}{2}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}$
$\quad+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\eta^{k}}{2\varsigma_{i}^{k}}d_{i}}\|\lambda^{k+1}-\lambda^{k}\|^{2}\Big),$
@@
Given some arbitrary initial points $\mathbf{x}^{0}\in\mathop{\bf dom}\phi$, $\theta^{0}\in\mathcal{K}^{*}$, $\lambda^{0}\in\mathbb{R}^{n|\mathcal{E}|}$, step sizes $\gamma^{k}>0$, $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$, and the momentum parameter $\eta^{k}>0$ for all $k\geq 0$, set $\mathbf{x}^{-1}=\mathbf{x}^{0}$, $\theta^{-1}=\theta^{0}$, $\lambda^{-1}=\lambda^{0}$, and consider the iterations as in (3.8). Suppose that there exists $\{t_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$ such that the algorithm parameters satisfy the following inequalities:
@@
Next, we argue that under the parameter rule in (4.13), for $k=1$ to K2K-2, we have $t_{k+1}Q^{k+1}(\mathbf{z})-t_{k}R^{k+1}(\mathbf{z})\leq 0$; indeed, it is easy to verify this inequality after writing it equivalently as follows, using the definitions of $Q^{k}(\mathbf{z})$ and $R^{k+1}(\mathbf{z})$ given in (4.5a) and (4.5b), respectively:
@@
Give arbitrary $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}\in\mathop{\bf dom}\phi$ and $\theta^{\prime},\theta^{\prime\prime}\in\mathcal{K}^{*}\cap\mathcal{B}$. Recall that $\Phi(\mathbf{x},\theta,\lambda)=\langle G(\mathbf{x}),\theta\rangle+\langle A\mathbf{x},\lambda\rangle$ for any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$. Thus, for any $i\in\mathcal{N}$ and $\lambda$, we have